OpenStudy (anonymous):

An algebraic expression has parentheses with unlike terms inside. Explain how to simplify this algebraic expression if a minus sign is in front of the parentheses.

OpenStudy (anonymous):

Distributive property: 2(x + y) = 2x + 2y -2(x + y) = -2x + -2y 1(x + y) = x + y Well, what's -1(x + y)? -x + -y Since anything multiplied by 1 is itself, -(x + y) = -1(x + y) = -x + -y = -x - y